OpenStudy (anonymous):

I'm looking out for maths tricks, specializing in "Square of a number" Like how to find square of numbers after 30 within seconds. There must be some tricks. Because the multiplying takes time, and it's hectic, and yep, without using calculators! Squares of big numbers! like numbers after 30. Or any random number that appears.

OpenStudy (anonymous):

There's a TED talk about a guy that can do this really fast.

OpenStudy (anonymous):

TED stands for?

OpenStudy (anonymous):

So there ain't no tricks?

OpenStudy (anonymous):

It's a site where people talk about interesting Ideas. Here's the link: http://www.ted.com/talks/lang/en/arthur_benjamin_does_mathemagic.html

OpenStudy (anonymous):

Don't know if it'll help you though.

OpenStudy (anonymous):

No worries, thank you for the link Thomas. The question is still open guys! :]

OpenStudy (kinggeorge):

It might be helpful to memorize the squares of 10, 20, 30, 40, 50, 60, 70, 80, 90, 100, 200, 300 up to some point. Fortunately, these are relatively easy to memorize.

OpenStudy (kinggeorge):

Also, if you can get the rules down for polynomial expansion, that could be very useful. E.g., \[312^2=(300+10+2)(300+10+2)\]\[312^2=300^2+10^2+2^2+2\cdot(300\cdot10+300\cdot2+10\cdot2)\]\[312^2=90000+100+4+2\cdot(3000+600+20)\]\[312^2=90104+2\cdot3620\]\[312^2=90104+7240\]\[312^2=97344\]

OpenStudy (kinggeorge):

Other than that, memorization and practice are the best way to go.